### v1

# Quantum Algorithm Benchmarks for PDEs Part 1: Evaluating Variational Quantum Eigensolver and Quantum Dynamics Algorithms on the Advection-Diffusion Equation

Abstract
We investigate the potential of near-term quantum algorithms in partial differential equations. Part 1 of this investigation focuses on solving a linear, one-dimensional PDE, namely the advection-diffusion equation, a convenient testbed to measure current capabilities of quantum computers. To the best of our knowledge, this work presents the first benchmarking study comparing a ground-state algorithm, specifically Variational Quantum Eigensolver (VQE), with three leading quantum dynamics algorithms, Trotterization, Variational Quantum Imaginary Time Evolution (VarQTE), and Adaptive Variational Quantum Dynamics Simulation (AVQDS), applied to the one-dimensional advection–diffusion equation on a small quantum computer. While Trotterization is a purely quantum approach, variational time-evolution algorithms (VarQTE, AVQDS) allow circuit-depth reductions suitable for NISQ devices. However, actual hardware results for these quantum dynamics methods incur sizable errors due to noise and a lack of systematic shot-noise statistics. To provide a noise-free baseline for algorithmic performance, we solve the advection-diffusion PDE using a VQE-based approach on a noiseless statevector simulator. Numerical experiments show that the VQE solution can achieve final-time infidelities as low as $\mathcal{O}(10^{-9})$ for $N=4$ qubits and moderate circuit depths. This compares favorably to the shot-based quantum dynamics approaches, whose hardware implementations show typical infidelities $\gtrsim\!10^{-2}$. By comparing noiseless VQE simulations to shot-based and hardware-deployed quantum dynamics methods, we assess both algorithmic accuracy and resource efficiency, providing a baseline for future performance evaluations on real quantum devices. We discuss limitations and future work, including exploring noise models and scaling up to multi-dimensional nonlinear PDEs relevant to engineering and financial applications.
I Introduction
Quantum computing (QC) has emerged as a promising paradigm to speed up the solution of high-dimensional partial differential equations (PDEs), which arise ubiquitously across fluid dynamics, finance, materials science, and machine learning [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Classical methods often suffer from exponential scaling in memory or runtime, especially for fine-resolution simulations. Recent advances in quantum algorithms, particularly variational quantum algorithms (VQAs) and quantum dynamics solvers, offer new tools for addressing these challenges on noisy intermediate-scale quantum (NISQ) hardware. Recent studies have begun to formalize quantum benchmarking approaches and metrics in the presence of noise, fostering a more systematic evaluation of quantum algorithms in near-term hardware environments [12, 13, 14].
One especially challenging domain is the numerical solution of partial differential equations (PDEs) describing transport phenomena (e.g., in fluid dynamics, heat transfer, and finance). Classical approaches often face high memory and runtime costs, particularly in higher dimensions or when fine mesh resolutions are required [15, 16]. Quantum algorithms, in principle, may alleviate these overheads by encoding the solution into amplitudes of a quantum state whose dimension grows exponentially with the number of qubits [17, 18, 19, 20, 21, 22].
Before fault-tolerant quantum hardware becomes available, one may still explore quantum algorithms for solving PDEs on today’s noisy intermediate-scale quantum (NISQ) devices [23, 24, 25, 26]. Variational quantum algorithms (VQAs) have emerged as a leading candidate for NISQ-era simulations due to their potential resilience against noise [27, 28, 29]. These hybrid quantum-classical schemes adopt parameterized circuits and iteratively update the parameters to minimize a suitable cost function, enabling approximate solutions to quantum chemistry, condensed matter, and PDE-based problems on small noisy hardware [19, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Compared to direct Hamiltonian simulation methods like Trotterization for qubits and qudits [58, 59, 60, 61], VQAs typically require fewer gates, although they involve classical optimization loops prone to problems such as barren plateaus [62, 63, 64, 65].
Recently, Alipanah et al. reported a pioneering attempt at solving the one-dimensional advection-diffusion equation on IBM quantum hardware via three quantum-dynamics algorithms (Trotterization, VarQTE, AVQDS) [66]. On a noiseless shot-based emulator, they achieved final infidelities on the order of $10^{-5}$, demonstrating the algorithms’ intrinsic capabilities when measurement outcomes are sampled but no hardware noise is present. However, in actual runs on a physical quantum processor, those infidelities increased to $\gtrsim 10^{-2}$, underscoring the impact of real hardware noise.
In addition to their hardware-based 1D studies, Alipanah et al. also explored a 2D advection-diffusion system at the simulator level. Although they showed promising results in a noise-free environment, the circuit sizes remain beyond the capabilities of current NISQ processors, making an actual hardware demonstration infeasible at present. This further underscores the importance of establishing noise-free benchmarks, such as those provided by idealized statevector simulations, to gauge the intrinsic performance of quantum PDE solvers before pushing toward higher-dimensional cases on real devices.
We benchmark a Variational Quantum Eigensolver (VQE) approach for the same PDE, but in an idealized (noise-free) statevector regime. By comparing to Alipanah et al.’s results, we can disentangle purely algorithmic performance (i.e., how well the PDE can be solved by an ansatz-based method) from hardware-induced errors. Our main contributions are presenting a VQE formulation of the PDE time-stepping by encoding each step into a ground-state problem of a suitable Hamiltonian, demonstrating on 4-qubit statevector simulations that VQE can achieve final-time infidelities $\approx 10^{-9}$ with relatively shallow circuits, far surpassing the typical $\gtrsim 10^{-2}$ errors reported on real hardware by Alipanah et al., and analyzing circuit resource usage.
Although the VQE approach in this paper is not immediately hardware-compatible in its current form (we do not measure partial gradients via shots), it provides a clean baseline for future shot-based or hardware implementations. This helps identify the true algorithmic potential, clarifies sources of error, and points to possible next steps in bridging the gap to real NISQ performance.
The rest of the paper is organized as follows. Section II reviews the one-dimensional advection-diffusion PDE and discretizes it via finite differences. Section III summarizes classical direct numerical simulations (DNS) and outlines how the PDE can be solved by either (i) the VQE-based linear-systems approach or (ii) quantum dynamics schemes (Trotter, VarQTE, AVQDS). Section III then presents our VQE results, including comparisons of gate counts to Alipanah et al., infidelities vs. time, and the effect of circuit depth. Finally, Section V concludes with an outlook on scaling to larger PDEs, the importance of noise analysis, and applications in finance and fluid engineering.
II Advection-Diffusion PDE and Quantum Encoding
We consider the dimensionless 1D advection-diffusion equation:
$\frac{\partial C}{\partial t}+\frac{\partial C}{\partial x}\;=\;\frac{1}{Pe}\,\frac{\partial^{2}C}{\partial x^{2}},$ (1)
with Péclet number $Pe=\tfrac{LU}{\Gamma}$ quantifying advective vs. diffusive transport. Here $C(x,t)$ is a scalar field, $0\leq x\leq 1$ is the domain (normalized by length $L$), and $t\geq 0$. Periodic boundary conditions are assumed. A second-order central finite difference scheme is applied in space:
$\frac{\partial C_{i}}{\partial t}+\frac{C_{i+1}-C_{i-1}}{2\,\Delta x}$ $=\frac{1}{Pe}\,\frac{C_{i+1}-2\,C_{i}+C_{i-1}}{\Delta x^{2}},$ (2)
with $i=0,\dots,2^{N}-1$ and $\Delta x=\tfrac{1}{2^{N}}$. We collect $C(x_{i})$ into a $2^{N}$-component vector $|C\rangle$ for quantum encoding, so that the $i$-th amplitude corresponds to $C(x_{i})$. For $N=4$, we have $16$ spatial points.
Rewriting (2) in vector form,
$\frac{\partial|C\rangle}{\partial t}\;=\;\hat{A}\,|C\rangle,$ (3)
$\hat{A}\equiv\frac{1}{Pe\,\Delta x}\begin{pmatrix}b&c&0&0&0&\dots&0&d\\ d&b&c&0&0&\dots&0&0\\ 0&d&b&c&0&\dots&0&0\\ 0&0&d&b&c&\dots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&&\vdots&\vdots\\ c&0&0&0&0&\dots&d&b\end{pmatrix},$ (4)
where $\quad b=-\frac{2}{\Delta x},\quad c=\frac{1}{\Delta x}-\frac{Pe}{2},\quad d=\frac{1}{\Delta x}+\frac{Pe}{2}$, and $\hat{A}$ is a $2^{N}\times 2^{N}$ operator (non-Hermitian, due to advection). We assume a normalized state vector $|C\rangle$ with periodic boundary conditions.
III Methodology: Classical (DNS), VQE, and Quantum Dynamics Approaches
Figure 1: (a) Scalar field $C(x)$ reconstructed from VQE outputs at various time points ($t=0$, $0.25$, $0.75$, $1.0$), alongside the corresponding DNS result at $t=1.0$ (dashed black line). (b) Infidelity $1-f$ between the VQE and DNS solutions vs. time for different circuit layers (1 to 10). At 5 layers, $1-f$ drops below $10^{-5}$. Beyond 5 layers, it can reach $10^{-9}$. (c) Final-time infidelity ($t=1$) vs. number of layers.
III.1 Classical DNS
A simple forward-Euler scheme updates $|C_{k}\rangle\mapsto|C_{k+1}\rangle$ in time:
$|C_{k+1}\rangle\;=\;\bigl{(}I+\Delta t\,\hat{A}\bigr{)}\,|C_{k}\rangle,$ (5)
for $k=0,1,\dots,N_{\mathrm{steps}}-1$, with time step $\Delta t=t_{\max}/N_{\mathrm{steps}}$. This direct numerical simulation (DNS) provides a stable reference solution, provided $\Delta t$ and $Pe$ satisfy stability constraints. In the results below, we label this classical vector as $|C_{\mathrm{DNS}}(t)\rangle$ and use it as ground truth to measure infidelities.
III.2 VQE-Based Linear-System Formulation
To implement the time-stepping (5) on a quantum device, we note that each step can be seen as solving a linear system:
$\bigl{[}I\;|\;-(I+\Delta t\,\hat{A})\bigr{]}\;\begin{pmatrix}|C_{k}\rangle\\ |C_{k+1}\rangle\end{pmatrix}\;=\;\begin{pmatrix}|C_{k}\rangle\\ 0\end{pmatrix}.$ (6)
One can rearrange this block-structured system of size $2^{N+1}\times 2^{N+1}$ to treat $|x\rangle=\begin{pmatrix}|C_{k}\rangle\\ |C_{k+1}\rangle\end{pmatrix}$ as the unknown. The unique solution vector that satisfies
$\mathcal{A}\,|x\rangle=\;|\mathbf{b}\rangle,\quad\mathcal{A}\in\mathbb{C}^{2^{N+1}\times 2^{N+1}},$
can be found by identifying the ground state of
$H_{k}\;=\;\mathcal{A}^{\dagger}\,\Bigl{(}I-|\mathbf{b_{k}}\rangle\langle\mathbf{b_{k}}|\Bigr{)}\,\mathcal{A}.$
Indeed, $H_{k}\,|x\rangle=0$ has the unique normalized solution $|x\rangle$ that encodes $|C_{k}\rangle$ in its first $2^{N}$ components and $|C_{k+1}\rangle$ in its last $2^{N}$ components. We then apply a VQE (variational quantum eigensolver) [19, 32] on $H_{k}$ to find this ground state. Repeating for $k=0,\dots,N_{\mathrm{steps}}-1$ yields the entire time evolution.
III.3 Quantum Dynamics Approaches: Trotter, VarQTE, AVQDS
An alternative is to directly simulate
$\frac{d}{d\beta}\,|C\rangle(\beta)=-\,i\,\hat{A}\,|C\rangle(\beta),\quad\beta=i\,t,$ (7)
which recasts the PDE in a Schrödinger-like equation with (complex) $\beta$. Trotterization decomposes $e^{-i\,\beta\hat{A}}$ into exponentials of Pauli strings, typically requiring many gates [67, 68]. Variational time-evolution methods (varQTE, AVQDS) replace the full operator with a parameterized ansatz and update angles in time [69, 70, 71], significantly lowering the gate count. Alipanah et al. implemented these algorithms in [66], reporting that Trotter circuits for $N=4$ can require $\sim 10^{5}$ entangling gates, whereas varQTE and AVQDS require $\sim 10^{2}$ gates but still suffer from hardware noise.
IV VQE Simulation Results
We solve the advection-diffusion equation (1) with $Pe=32$, domain $x\in[0,1]$, $N=4$ qubits, and a trapezoid-like initial condition. We choose a time step $\Delta t=0.002$ and run until $t=1.0$. At each step, we construct $H_{k}$ and run VQE on a statevector simulator (no shots). The circuit ansatz is a real-amplitudes (Ry+CNOT) design with variable depth (layers).
We employ a RealAmplitudes ansatz consisting of single-qubit $R_{y}(\theta)$ rotations and entangling gates (CNOTs) between neighboring qubits. The RealAmplitudes ansatz provides gate efficiency for PDEs involving real-valued states. It restricts the parameterized single-qubit gates to real-valued rotations, specifically using only $R_{y}$ gates. Figure 2 shows a schematic of the $N=4$ version of this ansatz, as well as a sample transpilation for an IBM-like linear qubit layout. Increasing the number of layers in this ansatz increases its expressivity but also its depth and parameter count. The optimizer adjust the parameters $\{\theta\}$ at each time step, aiming to minimize the VQE cost function that encodes our linear system. In the following, we report the infidelities and resource estimates of this approach, and compare them with those from other quantum dynamics algorithms.
Figure 2: Schematic of our real-amplitudes VQE ansatz for $N=4$ qubits, alongside its transpilation for a linear qubit topology. Each layer has single-qubit $R_{y}(\theta)$ gates followed by CNOTs between adjacent qubits.
Figure 1(a) shows the scalar field $C(x)$ reconstructed at times $t=0,0.25,0.75,1.0$ from the VQE amplitudes, for a circuit depth of 10 layers. The solution matches well with the DNS ground truth. Figure 1(b) plots the infidelity $1-f(t)$ where $f(t)=\lvert\langle C_{\mathrm{DNS}}(t)\mid C_{\mathrm{VQE}}(t)\rangle\rvert^{2}$. Even for as few as 5 layers, the final infidelity is below $10^{-5}$, and at 10 layers it reaches $\approx 10^{-9}$. Figure 1(c) further illustrates that increasing the number of layers improves final-time fidelity dramatically. In contrast, the shot-based real hardware runs of varQTE/AVQDS in [66] typically showed final infidelities in the range of $10^{-2}$$10^{-1}$ when $N=4$ and $t=1.0$, likely dominated by hardware noise and shot uncertainty. Our statevector-based VQE thus sets an ideal baseline for how accurately the PDE could be solved with a modest-depth circuit in a noise-free setting.
Table 1 compares single-/two-qubit gate counts for Trotterization, varQTE, AVQDS (as reported by [66]), and our VQE approach with 5 layers (similar to the preceding figure). The Trotter approach demands $\sim 10^{5}$ total gates, overshadowing the $\sim 10^{2}$ gates of the variational algorithms. Among the latter, we see VQE has $\sim 15$ two-qubit (CNOT) gates, which is even fewer than AVQDS or varQTE. Our circuit depth also remains modest (55), benefiting from the real-amplitudes structure. In principle, VQE must be repeated at each time step, so the total resource cost scales with $N_{\mathrm{steps}}$. That said, for the small example here ($N=4$ and $500$ steps), the overall gate count is still well below Trotter’s $\gtrsim 10^{5}$ for a single step. For more advanced PDEs or 2D grids, further circuit optimizations (e.g., entanglement-preserving layout) or quantum data compression [72, 73] may become essential.
Table 1: Representative gate counts and circuit depths for $N=4$ qubits. Trotterization is from [66], requiring $\sim 10^{5}$ gates. VarQTE and AVQDS also from [66], each with $\mathcal{O}(10^{2})$ gates. Our VQE approach (5 layers) uses fewer total gates and only 15 two-qubit gates, leading to a depth of 55.
Gate Trotter VarQTE AVQDS VQE
$X$ 317 0 6 0
$\sqrt{X}$ 53646 108 79 78
$RZ$ 48460 109 67 93
cz 20213 30 40 15
Total 122636 247 192 186
Depth 76021 90 129 55
V Conclusion and Outlook
We have benchmarked a Variational Quantum Eigensolver (VQE) approach to solving the 1D advection-diffusion PDE, comparing it to the Trotter and variational real and imaginary time-evolution methods of Alipanah et al. [66]. Whereas those methods were implemented on IBM hardware and subject to significant noise, we used an idealized statevector simulator to assess the intrinsic performance of a VQE-based linear-systems solver. Numerical results for $N=4$ qubits demonstrate final-time infidelities as small as $10^{-9}$ with only tens of two-qubit gates per time step.
In this paper, we focus on a statevector simulation of a VQE-based approach, providing a noise-free baseline for algorithmic performance. While our results are comparable in spirit to Alipanah et al.’s noiseless emulator outcomes, they differ in two key respects. First, our VQE approach bypasses any sampling error by computing exact expectation values from the full wavefunction, whereas Alipanah et al.’s shot-based emulator still includes measurement statistics (albeit with no hardware noise). Thus, if repeated measurement outcomes were simulated, we might anticipate slightly higher infidelities than in the ideal statevector model. Second, we solve the PDE via a VQE linear-system formulation rather than simulating real or imaginary-time evolution. Both approaches are valid but differ in circuit structure and required gates. Hence, direct comparisons in overall gate counts or final errors can be nuanced.
Nevertheless, comparing our noise-free (statevector) VQE results to the noise-free shot-based emulator—and ultimately to their noisy hardware runs—helps disentangle purely algorithmic performance from the hardware’s limitations. Our final infidelities, as low as $10^{-9}$ in the idealized scenario, illustrate the potential headroom before noise becomes dominant. Moving forward, exploring shot noise and explicit noise models (e.g. amplitude damping or depolarizing channels) will bridge the gap between idealized results and the realities of NISQ hardware performance.
Several key takeaways emerge. In a noise-free setting, VQE can converge extremely close to the PDE solution even with moderate circuit depth. This illustrates the potential power of variational PDE solvers when hardware error is not the limiting factor, consistent with recent VQE benchmarks for spin models on near-term hardware [54]. Trotterization is still gate-intensive ($\sim 10^{5}$) for $N=4$. By contrast, VQE is comparable in gate count to varQTE or AVQDS yet obtains robust solutions. This suggests that, for linear PDEs, VQE-based approaches can be competitive in the near term. Alipanah et al. did not isolate or statistically characterize shot noise, but their final solutions were degraded by hardware errors. Our statevector results highlight how far the hardware is from the algorithmic ideal.
In future work, we plan to incorporate hardware-like noise models (e.g., amplitude damping, depolarizing channels, readout error) into VQE PDE simulations and measure the effect on fidelity. We will also investigate scaling to two-dimensional PDEs relevant for incompressible flows and PDEs from finance (e.g., [74, 75, 76]). The persistent challenge is to maintain accuracy while controlling circuit depth and mitigating NISQ noise, possibly through error mitigation [77, 78], dynamical decoupling [79, 80], or specialized ansätze [81, 82].
Overall, our results indicate that near-term quantum hardware will continue to face noise-related limitations, but from an algorithmic standpoint, VQE is a viable approach for small PDE grids, on par with or better than existing quantum-dynamics methods. As hardware matures, combining VQE-like strategies with advanced error mitigation, data compression, and adaptive circuit design could open a path toward quantum advantage in solving PDEs at scale.
Acknowledgments
The author acknowledges discussions with his spouse. The author acknowledges discussions with the authors of the concurrent work [66] and the support from the University of Pittsburgh and Pittsburgh Quantum Institute. The author acknowledges valuable discussions at the 2024 Quantum Industry Day, held at the Simons Institute for the Theory of Computing in Berkeley, California, and the AI+Quantum Workshop at the Aspen Center for Physics in February 2025. This work was supported in part by the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452.